OpenStudy (anonymous):

solve the system of equations by graphing x-y=-3 x+y=5

OpenStudy (anonymous):

(1, 4)

OpenStudy (anonymous):

Pls can u show me how u arrived at that

OpenStudy (agreene):

http://www.wolframalpha.com/input/?i=plot+%7Bx-y%3D-3+and+x%2By%3D5%7D via graphing, as they requested.

OpenStudy (anonymous):

First I rewrote the equations in y= mx + b format. x - y = -3 -y = -x-3 y = x + 3 x + y = 5 y = -x+5 Then I used the slope and y-intercepts to plot both equations on a graph. The point they intersected on was (1, 4).

OpenStudy (anonymous):

Thank u Emmy82650

OpenStudy (anonymous):

You're welcome!